OpenStudy (anonymous):

How would I find the exact value of sin(7pi/8) using a half-angle identity?

OpenStudy (mertsj):

7pi/8 is half of 14pi/8 which is 7pi/4. So use the half angle formula and use 7pi/4 for x

OpenStudy (anonymous):

How can you change it like that?, because 7pi/8 isn't equal to 7pi/4.. That confuses the crud out of me.

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ \cfrac{7}{8}\cdot 2\implies \cfrac{7}{4}\qquad \cfrac{7\pi}{8}\cdot 2\implies \cfrac{7\pi}{4}}}\\ \quad \\ sin\left(\frac{7\pi}{8}\right)\implies sin\left(\cfrac{\frac{7\pi}{4}}{2}\right)\)

OpenStudy (anonymous):

Oh, okay. So the only reason he could do that was because the problem has a two in it?

OpenStudy (jdoe0001):

yes \(\bf sin\left(\frac{7\pi}{8}\right)\implies sin\left(\cfrac{\frac{7\pi}{4}}{2}\right)=\pm\sqrt{\cfrac{1-cos\left(\frac{7\pi}{4}\right)}{2}}\)

OpenStudy (anonymous):

Are my answers supposed to be deciamls, or the ones with pi in them? I have no idea how to do this...

OpenStudy (jdoe0001):

well.. what's the \(\bf cos\left(\frac{7\pi}{4}\right)\quad ?\)

OpenStudy (anonymous):

.707?

OpenStudy (jdoe0001):

well.. or one can say \(\bf \cfrac{\sqrt{2}}{2}\) check your Unit Circle, thus \(\bf sin\left(\cfrac{\frac{7\pi}{4}}{2}\right)=\pm\sqrt{\cfrac{1-cos\left(\frac{7\pi}{4}\right)}{2}} \implies sin\left(\cfrac{\frac{7\pi}{4}}{2}\right)=\pm\sqrt{\cfrac{1-\frac{\sqrt{2}}{2}}{2}} \\ \quad \\ sin\left(\cfrac{\frac{7\pi}{4}}{2}\right)=\pm\sqrt{\cfrac{\frac{\sqrt{2-\sqrt{2}}}{2}}{2}}\implies sin\left(\cfrac{\frac{7\pi}{4}}{2}\right)=\pm\sqrt{2-\sqrt{2}}\)

OpenStudy (anonymous):

I can't see it...

OpenStudy (jdoe0001):

?

OpenStudy (jdoe0001):

refresh the page maybe

OpenStudy (anonymous):

It's not that, it goes off the right side too far.

OpenStudy (jdoe0001):

hmm ok... \(\bf sin\left(\cfrac{\frac{7\pi}{4}}{2}\right)=\pm\sqrt{\cfrac{1-cos\left(\frac{7\pi}{4}\right)}{2}} \\ \quad \\\implies sin\left(\cfrac{\frac{7\pi}{4}}{2}\right)=\pm\sqrt{\cfrac{1-\frac{\sqrt{2}}{2}}{2}} \\ \quad \\ sin\left(\cfrac{\frac{7\pi}{4}}{2}\right)=\pm\sqrt{\cfrac{\frac{\sqrt{2-\sqrt{2}}}{2}}{2}}\\ \quad \\\implies sin\left(\cfrac{\frac{7\pi}{4}}{2}\right)=\pm\sqrt{2-\sqrt{2}}\)

OpenStudy (anonymous):

Haha. I just did the exact same thing! :p Okay, two questions. #1) How did you get sqrt2/2? And #2) What is that third step? How did we go from 1-sqrt2/2/2, to sqrt2-sqrt2/2/2???

OpenStudy (jdoe0001):

#1) by using the Unit Circle :), I could have gotten it from the pythagorean theorem anyway #2) is a fractions addition, so \(\bf 1-\cfrac{\sqrt{2}}{2}\implies \cfrac{1}{1}-\cfrac{\sqrt{2}}{2}\implies \cfrac{2-\sqrt{2}}{2} \)

OpenStudy (jdoe0001):

I'm a bit lagged

OpenStudy (anonymous):

Oh, okay, I get the first part. I still don't get the second. Maybe I need to go back to middle school. How did the top 1 become a 2?

OpenStudy (jdoe0001):

well... that's just a fractions' addition.. you use the LCD for that

OpenStudy (anonymous):

Me too. Takes it a million years to catch up on my typing..

OpenStudy (anonymous):

Oh. Got it. I promise I am nowhere near as dumb as I seem right now.

OpenStudy (anonymous):

How did you use the Unit Circle to get sqrt2/2?

OpenStudy (jdoe0001):

http://www.sciencedigest.org/UnitCircle.gif